OpenStudy (jenniferjuice):

What is the axis of symmetry for f(x) = −3x2 + 12x − 6? x = 4 x = −4 x = 2 x = −2

OpenStudy (anonymous):

What is x-coordinate of the vertex ?

OpenStudy (jenniferjuice):

-3 ? idk

OpenStudy (anonymous):

-b/2a

OpenStudy (anonymous):

=?

OpenStudy (anonymous):

f(x) = ax^2 +bx+c, a =? b=? c=?

OpenStudy (jenniferjuice):

?what do i do?

OpenStudy (anonymous):

Solve for x: -(3 x^2)+12 x-6 = 0 Write the quadratic equation in standard form. Divide both sides by -3: x^2-4 x+2 = 0 Solve the quadratic equation by completing the square. Subtract 2 from both sides: x^2-4 x = -2 Take one half of the coefficient of x and square it, then add it to both sides. Add 4 to both sides: x^2-4 x+4 = 2 Factor the left hand side. Write the left hand side as a square: (x-2)^2 = 2 Eliminate the exponent on the left hand side. Take the square root of both sides: x-2 = sqrt(2) or x-2 = -sqrt(2) Look at the first equation: Solve for x. Add 2 to both sides: x = 2+sqrt(2) or x-2 = -sqrt(2) Look at the second equation: Solve for x. Add 2 to both sides: Answer would be | x = 2+sqrt(2) or x = 2-sqrt(2)

OpenStudy (anonymous):

oh boy!! \[f(x) = \color{red}{a}x^2 +\color{blue}{b}x + c\\f(x) = \color{red}{-3}x^2 +\color{blue}{12}x + c\] tell me a=?? b=??

OpenStudy (anonymous):

@jenniferjuice I guide you step by step to solve the problem. You have to show me your trying. If you don't want to study, It's perfectly ok to me. :) take your time to answer me or you can keep silent.

OpenStudy (jenniferjuice):

sorry i was away from the computer excuse me

OpenStudy (anonymous):

ok, still need know how to get?? if so, answer me, a =? b=?

OpenStudy (jenniferjuice):

a=-3 b=12

OpenStudy (anonymous):

yup, what is \(\dfrac{-b}{2a}\) ? which is x-coordinate of the vertex.

OpenStudy (jenniferjuice):

-12/2*-3

OpenStudy (jenniferjuice):

=2

OpenStudy (anonymous):

yup, the line passes through the x-coordinate of the vertex is the symmetric line of the graph.